# SMART: Submodular Data Mixture Strategy for Instruction Tuning

### v2

@@ 1 Introduction
  • (RQ1) Given a huge multitask instruction-tuning dataset and a limited fine-tuning budget which is defined by the total number of $(prompt,response)$ pairs that can be used for fine-tuning, how do we divide this budget among thousands of tasks present in the dataset? i.e., how many instances to sample from each task? and which instances to sample from each task?
  • (RQ2) Can we go a step further and strategically prune some tasks altogether and only finetune on a small subset of representative tasks without hurting the performance? If yes, what is the nature of this subset?
@@
As a first step towards addressing both of the above RQs, we first define a common subset selection problem (more formally stated in Section 3) as follows - Given a huge collection of $M$ instruction-formatted task datasets, a task budget $M^{\prime}\leq M$ and a total budget $N^{\prime}$, of $(prompt,response)$ pairs, which $M^{\prime}$ tasks to select, how many instances to select from each of these $M^{\prime}$ tasks and which instances to select? Note that RQ1 is an instance of this problem where $M^{\prime}=M$.
@@
  • We introduce SMART — a novel data mixture strategy for instruction tuning that models the data mixture problem (defined in Section 3) as a sequence of two cardinality-constrained submodular maximization problems (Section 3) and offer empirical evidence that it outperforms both examples proportional and equal mixing baselines (Section 4).
  • Existing works (Iyer et al., 2022; Longpre et al., 2023) have reported a continuous increase in performance upon increasing the number of tasks (though the gains themselves may be diminishing). However, we posit that this depends on the order in which new tasks are incorporated and show empirically that in the case of SMART mixtures, a performance peak is observed with an initial addition of few representative tasks and upon adding more tasks, the performance is not sustained (Section  4.6).
@@ 2 Background: Submodularity
The above cardinality-constrained submodular maximization problem is NP-complete Feige (1998). However, if $f$ is monotone submodular (i.e., $f(X)\leq f(Y)$ whenever $X\subseteq Y$), Nemhauser et al., 1978; Fisher et al., 1978 show that a simple greedy algorithm described in Algorithm 1 can be used to find an approximate solution ($\mathcal{S}^{greedy}$) with a guarantee that $f(\mathcal{S}^{greedy})\geq(1-1/e)f(\mathcal{S}^{*})$. [^1] Assuming $P\neq NP$, this is the best approximation ratio that can be achieved by any polynomial time algorithm.
@@ 3 Approach
SMART models the above problem as a sequence of two cardinality-constrained submodular maximization problems. The first one is to select a weighted subset of $M^{\prime}$ tasks and the second one is to select instances a total of $N^{\prime}$ instances from these tasks. The weights obtained in the first stage will be used to determine how many instances we sample from each task.
@@ 3.1.1 Stage-1: Weighted Task Subset Selection
In this first stage, given the instruction-tuning dataset $\mathcal{D}=\{\mathcal{T}_{1},\dots,\mathcal{T}_{M}\}$, our goal is to find $\mathcal{D}^{\prime}=\{\mathcal{T}_{1}^{\prime},\dots,\mathcal{T}_{M^{\prime}}^{\prime}\}$($\mathcal{D}^{\prime}\subseteq\mathcal{D}$), along with the instance budgets, $\{N_{1}^{\prime},\dots,N_{M^{\prime}}^{\prime}\}$, such that $\sum_{j=1}^{M^{\prime}}|N_{j}^{\prime}|=N^{\prime}$.
@@ 3.1.2 Stage-2: Instance Subset Selection
  $encoder\leftarrow SentenceEncoder()$
  $prompt\_embeddings\leftarrow dict()$
  $task\_embeddings\leftarrow dict()$
@@
  $\mathcal{K}_{task}=cos\_sim(task\_embeddings)$
@@
     $\mathcal{K}_{\mathcal{T}_{j}^{\prime}}=cos\_sim(prompt\_embeddings[\mathcal{T}_{j}^{\prime}])$
@@ 3.2 Obtaining Similarity Measures
In Table 1, the three submodular functions require similarity measures ($s_{ij}$) between items in the ground set. For instance subset selection, cosine similarity between prompt embeddings is used. For weighted task subset selection, cosine similarity between task embeddings is calculated, where the task embeddings are computed as the average prompt embeddings, following Vu et al. (2020). Although there are other sophisticated methods  (Achille et al., 2019; Zhou et al., 2022; Xi et al., 2023; Vu et al., 2021) for obtaining task embeddings, most of them depend on the LLM at hand.
@@ 3.3 Choosing $f_{1}$ and $f_{2}$
Each submodular function in Table 1 serves a different purpose: facility location emphasizes representation, graph cut balances representation and diversity, and logDeterminant prioritizes diversity. We treat $f_{1}$ and $f_{2}$ as hyperparameters in our grid search in Section 4.7, exploring three options for each function. Determining the optimal $f_{1}$ and $f_{2}$ is part of our research question (RQ2). In Section 4.9, we discuss the grid search findings and their qualitative implications for instruction tuning.
@@ 4.1 Finetuning Data
We encode prompts in the collection with GTE-large (Li et al., 2023b), a light-weight (340M parameters) BERT-based effective  (Muennighoff et al., 2022) sentence encoder for semantic textual similarity. Task embeddings are obtained by averaging the prompt embeddings (Section 3.2). We use SUBMODLIB (Kaushal et al., 2022), which has the necessary algorithms implemented, to obtain the weighted task subsets (Section  3.1.1) and the instances (Section  3.1.2). We distribute the instance budgets equally among task templates based on findings by Longpre et al. (2023) which showed that an equal number of zero-shot and few-shot templates yield the best performance on held-out tasks.
@@ 4.2 Finetuning Procedure
We evaluate SMART on three 7B parameter LLMs: Llama-2 Touvron et al. (2023), Falcon Almazrouei et al. (2023), and Mistral Jiang et al. (2023). We fine-tune each model for 1 epoch on a given data mixture with a learning rate of $52e-$ 5 for Llama-2 and Falcon, and $65e-$ 6 for Mistral. A batch size of $64$, weight decay of $0.1$, and cosine learning rate decay with a linear warmup for the initial $1\%$ training steps are employed. The experiments run on 8 NVIDIA A100-SXM4-80GB GPUs, utilizing Flash Attention Dao (2023) for memory efficiency and speeding up finetuning. Our code is open-sourced here.
@@ 4.3 Baselines
We compare SMART with three baseline mixture strategies: Examples Proportional Mixture, Equal Mixture  (Raffel et al., 2020) and FLANv2 Mixture (Longpre et al., 2023)
@@
FLANv2 Mixture (Baseline-3)
Following Longpre et al. (2023), 40%, 32%, 20%, 5% and 3% are assigned to the 5 submixtures of FLAN 2022 - FLANv1, T0, NIV2, CoT and Dialog respectively and instances are sampled randomly from each submixture.
@@
Table 2: Comparison of SMART with baselines on MMLU-zeroshot and BBH-zeroshot. EPM (Baseline-1) denotes Examples Proportional Mixture and EM (Baseline-2) denotes Equal Mixture. All the metrics are Exact Matches.
@@ 4.4 Evaluation Protocol
We evaluate the finetuned models on two benchmark datasets: MMLU Hendrycks et al. (2020) with 57 tasks assessing world knowledge and problem-solving, and BBH Suzgun et al. (2022) with 23 challenging tasks from Big-Bench Srivastava et al. (2022). MMLU covers STEM, Humanities, Social Sciences, and Other (business, medical, and misc.) categories, while BBH includes NLP and Algorithmic tasks [^2] The Algorithmic subcategory is so named because these tasks (e.g., 2-digit arithmetic) do not require an LLM to be solved.. Evaluation involves prompting the LLM directly and using Exact Match as the scoring metric. Responses undergo basic post-processing steps (removing punctuation and lowercasing) and are generated using Greedy Decoding. For baseline data mixtures, 3 mixtures with different random seeds are created and the mean exact match of the 3 finetuning runs is reported.
@@ 4.5 Addressing RQ1 ($M^{\prime}=M$)
RQ1 is an instance of the subset selection problem from Section  3, where $M^{\prime}=M$, allowing all tasks for finetuning; however we still have a constraint ($N^{\prime}$) on total number of $(prompt,response)$ pairs. We use Graph-Cut and Facility Location functions from Table 1 in Stage-1 (Section 3.1.1) and Stage-2 (Section 3.1.2) of SMART respectively. This choice of $f_{1}$ and $f_{2}$ is determined via grid search discussed in Section 4.7. Table 2 compares SMART with baselines (on evaluation benchmarks described in Section 4.4) by finetuning Llama-2-7b, on data mixtures generated by varying $N^{\prime}$ in $\{25000,50000,100000,200000,400000\}$. As can be inferred from the table, SMART consistently performs better than all the three baselines and also performs close to finetuning on full collection.
@@
Table 3: Grid Search on submodular functions $f_{1}$ and $f_{2}$. FL denotes Facility Location, GC denotes Graph Cut and LOGDET denotes Log-Determinant
4.6 Addressing RQ2 ($M^{\prime}<M$)
To study the effect of pruning some tasks altogether, on both model and the baselines (RQ2), we finetune on both the baseline and SMART data mixtures by varying the number of tasks ($M^{\prime}$) in $\{8,16,32,64,128,256,512,1024,1840\}$ and the total number of instances ($N^{\prime}$) in $\{25000,50000,100000,200000,400000\}$. Figure  2 shows task-scaling plots for each $N^{\prime}$. Baselines’ performance steadily improves with more tasks (even though the gains are diminishing after a point), but SMART shows an optimal performance at a certain in-between point which isn’t maintained consistently upon adding more tasks, suggesting that rather than increasing tasks, focusing on representative ones and sampling more instances from these might be more beneficial in low-budget scenarios. Even with an ample budget, scaling tasks should be judicious, as close to  97% of performance achievable by using the entire FLAN collection can be attained with a subset of 16 representative tasks alone with $N^{\prime}=200000$.
@@ 4.6 Addressing RQ2 ($M^{\prime}<M$)
Figure 2: Task-Scaling Curves for Llama-2 (Y-axis is the weighted average of Exact Matches on MMLU and BBH)
4.7 Ablation Study: Submodular Function
Both stages of SMART require selecting submodular functions ($f_{1},f_{2}$), treated as hyperparameters due to uncertainty about suitable functions for instruction tuning (Section 3.3). We conduct a grid search on $f_{1}$ and $f_{2}$ using three functions from Table 1, by setting $M^{\prime}=M$ and varying $N^{\prime}$ in $\{25000,50000,100000,200000\}$. In this work, we use $\lambda=0.4$ for Graph-Cut. Results in Table 3 indicate Graph Cut is optimal for task subset selection, while Facility Location is best for instance selection when $M^{\prime}=M$. The task scaling curves for each combination of $f_{1}$ and $f_{2}$ (for different $N^{\prime}$s) are in Appendix B. However, here we highlight the case where $f_{1}$ is Graph Cut. We find that optimal $f_{2}$ in this case also depends on $M^{\prime}$. For instance, in Figure 1 where $N^{\prime}=200000$, Facility Location performs best at higher $M^{\prime}$, indicating the importance of representation, while Graph Cut and Log Determinant show better performance at lower $M^{\prime}$, suggesting the importance of diversity. We hypothesize this is because for higher $M^{\prime}$s, each task on average gets a relatively low budget and the need for representation dominates the need for diversity; however when there is sufficient enough budget for each task, i.e., for lower $M^{\prime}$s, the need for diversity takes over.
@@ 4.8 Ablation Study: LLM
We also test the effectiveness of SMART strategy on two other LLMs - Falcon-7B and Mistral-7B. Table 4 presents a comparison of SMART mixture with the baseline mixtures for Falcon and Mistral when $M^{\prime}=M$. For Falcon, we only report MMLU-zero-shot since Falcon gets an exact match of 0.0 for BBH-zero-shot even after finetuning. The task-scaling curves for these models are present in Appendix C.

### v3

@@ 1 Introduction
  • (RQ1) Given a huge multitask instruction-tuning dataset and a limited fine-tuning budget which is defined by the total number of $(prompt,response)$ instances that can be used for fine-tuning, how do we divide this budget among thousands of tasks present in the dataset? i.e., how many instances to sample from each task? and which instances to sample from each task?
  • (RQ2) Can we go a step further and strategically prune some tasks altogether and only fine-tune on a small subset of representative tasks without hurting the performance? If yes, what is the nature of this subset?
@@
As a first step towards addressing both of the above RQs, we first define a common subset selection problem (more formally stated in Section 3) as follows - Given a huge collection of $M$ instruction-formatted task datasets, a task budget $M^{\prime}\leq M$ and a total budget ($N^{\prime}$) of $(prompt,response)$ pairs, which $M^{\prime}$ tasks to select? and how many instances to select from each of these $M^{\prime}$ tasks and which instances to select? Note that RQ1 is an instance of this problem where $M^{\prime}=M$.
@@
  • We introduce SMART — a novel data mixture strategy for instruction tuning that models the data mixture problem (Section 3) as a sequence of two cardinality-constrained submodular maximization problems and offer empirical evidence that it outperforms both examples proportional and equal mixing baselines (Section 4) as well as the mixture weights proposed by Longpre et al. (2023).
  • Existing works like Longpre et al. (2023) have reported a continuous increase in performance upon increasing the number of tasks (though the gains themselves may be diminishing). However, we posit that this depends on the order in which new tasks are incorporated and show empirically that in the case of SMART mixtures, a performance peak is observed with an initial addition of few representative tasks and upon adding more and more tasks, the performance is not sustained (Section  4.6).
@@ 2 Background: Submodularity
The above cardinality-constrained submodular maximization problem is NP-complete Feige (1998). However, if $f$ is monotone submodular (i.e., $f(X)\leq f(Y)$ whenever $X\subseteq Y$), Nemhauser et al., 1978; Fisher et al., 1978 show that a simple greedy algorithm described in Algorithm 1 can be used to find an approximate solution $\mathcal{S}^{greedy}$, with a guarantee that $f(\mathcal{S}^{greedy})\geq(1-1/e)f(\mathcal{S}^{*})$. [^1] Assuming $P\neq NP$, this is the best approximation ratio that can be achieved by any polynomial time algorithm.
@@ 3 Approach
SMART models the above problem as a sequence of two cardinality-constrained submodular maximization problems. The first one is to select a weighted subset of $M^{\prime}$ tasks and the second one is to select instances - a total of $N^{\prime}$ instances from these tasks. The weights obtained in the first stage will be used to determine how many instances we sample from each task.
@@ 3.1.1 Stage-1: Weighted Task Subset Selection
In this first stage, given the instruction-tuning dataset $\mathcal{D}=\{\mathcal{T}_{1},\dots,\mathcal{T}_{M}\}$, our goal is to find $\mathcal{D}^{\prime}=\{\mathcal{T}_{1}^{\prime},\dots,\mathcal{T}_{M^{\prime}}^{\prime}\}$ where $\mathcal{D}^{\prime}\subseteq\mathcal{D}$, along with the instance budgets, $\{N_{1}^{\prime},\dots,N_{M^{\prime}}^{\prime}\}$, such that $\sum_{j=1}^{M^{\prime}}|N_{j}^{\prime}|=N^{\prime}$.
@@ 3.1.2 Stage-2: Instance Subset Selection
  $encoder\leftarrow\text{SentenceEncoder()}$()
  $prompt\_embeddings\leftarrow\text{dict()}$()
  $task\_embeddings\leftarrow\text{dict()}$()
@@
  $\mathcal{K}_{task}=\text{cos\_sim}(task\_embeddings)$
@@
     $\mathcal{K}_{\mathcal{T}_{j}^{\prime}}=\text{cos\_sim}(prompt\_embeddings[\mathcal{T}_{j}^{\prime}])$
@@ 3.2 Obtaining Similarity Measures
The three submodular functions listed in Table 1 require computation of similarity measures ($s_{ij}$) between items in the ground set. So, all prompts in the dataset are first encoded using a sentence encoder. For instance subset selection (i.e., Stage-2), cosine similarity between prompt embeddings is used as the similarity measure and for weighted task subset selection (i.e., Stage-1), cosine similarity between task embeddings is computed, where the task embeddings are computed as the average prompt embeddings, following Vu et al. (2020). Although there are other sophisticated methods  (Achille et al., 2019; Zhou et al., 2022; Xi et al., 2023; Vu et al., 2021) for obtaining task embeddings, most of them depend on the LLM at hand.
@@ 3.3 Choosing $f_{1}$ and $f_{2}$
Each submodular function in Table 1 captures a different property: facility-location emphasizes representation, graph-cut balances representation and diversity, and log-determinant prioritizes diversity. We treat $f_{1}$ and $f_{2}$ as hyperparameters in our grid search in Section 4.8, exploring three options for each function. Determining the optimal $f_{1}$ and $f_{2}$ is part of our research question (RQ2). We discuss the findings of grid search and their qualitative implications for instruction tuning in Section 4.10.
@@ 4.1 Finetuning Data
We encode prompts in the collection with GTE-large (Li et al., 2023b), a light-weight (340M parameters) BERT-based effective  (Muennighoff et al., 2022) sentence encoder for semantic textual similarity. Task embeddings are obtained by averaging the corresponding prompt embeddings (Section 3.2). We use SUBMODLIB (Kaushal et al., 2022), which has the necessary algorithms implemented, to obtain the weighted task subsets (Section  3.1.1) and the instances (Section  3.1.2). We distribute the instance budgets equally among task templates based on findings by Longpre et al. (2023) which showed that an equal number of zero-shot and few-shot templates yield the best performance on held-out tasks.
@@ 4.2 Finetuning Procedure
We evaluate SMART on three 7B parameter LLMs: Llama-2 Touvron et al. (2023), Falcon Almazrouei et al. (2023), and Mistral Jiang et al. (2023). We fine-tune each model for 1 epoch on a given data mixture with a learning rate of $52e-$ 5 for Llama-2 and Falcon, and $65e-$ 6 for Mistral. A batch size of $64$, weight decay of $0.1$, and cosine learning rate decay with a linear warmup for the initial $1\%$ training steps are employed. The experiments all ran on 8 NVIDIA A100-SXM4-80GB GPUs, utilizing Flash Attention Dao (2023) for memory efficiency and speeding up finetuning. Our code is open-sourced here.
@@ 4.3 Baselines
We mainly compare SMART with two baseline mixture strategies: Examples Proportional Mixture, Equal Mixture  (Raffel et al., 2020).
@@
We also compare SMART with mixture weights used by Longpre et al. (2023) but since their mixture weights apply to the five sub mixtures listed in Section 4.1 rather than individual tasks, we analyse this separately in Section 4.7.
@@
Table 2: Comparison of SMART with baselines on MMLU-zeroshot and BBH-zeroshot for Llama-2-7b. EPM (Baseline-1) denotes Examples Proportional Mixture and EM (Baseline-2) denotes Equal Mixture. All the scores are Exact Matches. Weighted average on 57 MMLU tasks and 23 BBH tasks is reported in the last column. For baselines, exact matches are obtained by averaging across 3 fine-tuning runs.
@@ 4.4 Evaluation Protocol
We evaluate the fine-tuned models on two benchmark datasets: MMLU Hendrycks et al. (2020) with 57 tasks assessing world knowledge and problem-solving, and BBH Suzgun et al. (2022) with 23 challenging tasks from Big-Bench Srivastava et al. (2022). MMLU covers STEM, Humanities, Social Sciences, and Other (business, medical, and misc.) categories, while BBH includes both NLP and Algorithmic tasks [^2] The Algorithmic subcategory is so named because these tasks (e.g., 2-digit arithmetic) do not require an LLM to be solved.. Evaluation involves prompting the LLM directly and using Exact Match as the scoring metric. Responses are generated using the greedy decoding approach and they undergo basic post-processing steps (removing punctuation and lower-casing) before calculating exact match. For baseline data mixtures, 3 mixtures with different random seeds are created and the mean exact match of the 3 fine-tuning runs is reported.
@@ 4.5 Addressing RQ1 ($M^{\prime}=M$)
RQ1 is an instance of the subset selection problem defined in Section  3, where $M^{\prime}=M$, allowing all tasks for finetuning; however we still have a constraint ($N^{\prime}$) on total number of $(prompt,response)$ pairs. We use Graph-Cut and Facility Location functions (Table 1) in Stage-1 (Section 3.1.1) and Stage-2 (Section 3.1.2) of SMART respectively. This choice of $f_{1}$ and $f_{2}$ is determined via grid search discussed in Section 4.8. Table 2 contains comparison of SMART mixtures with baseline mixtures, on MMLU and BBH benchmarks, upon instruction fine-tuning Llama-2-7b on data mixtures generated by varying $N^{\prime}$ in $\{25000,50000,100000,200000,400000\}$. SMART data mixtures consistently perform better than both examples proportional mixtures and equal mixtures baseline.
Figure 1: Task-Scaling Curves in the case of Llama-2-7b for different $N^{\prime}$s, where X-axis is the number of tasks ($M^{\prime}$) and Y-axis is the weighted average of Exact Matches on MMLU and BBH.
4.6 Addressing RQ2 ($M^{\prime}<M$)
RQ2 is an instance of the subset selection problem defined in Section 3, where $M^{\prime}<M$ i.e., studying the effect of pruning some tasks altogether, on both SMART and the baseline data mixtures. We fine-tune Llama-2-7b on both the baseline and SMART data mixtures by varying the number of tasks ($M^{\prime}$) in $\{8,16,32,64,128,256,512,1024,1840\}$ and the total number of instances ($N^{\prime}$) in $\{25000,50000,100000,200000,400000\}$. Figure 1 depicts the task-scaling plots for each $N^{\prime}$. Baseline data mixtures’ performance steadily improves upon increasing the number of tasks, even though the gains are diminishing after a point. In contrast, SMART data mixtures yield optimal performance at an in-between point and this performance does not seem to sustain upon adding more and more tasks, suggesting that rather than increasing the tasks, focusing on representative tasks and sampling more instances from these might be more beneficial in low-budget scenarios. Even with an ample budget, scaling tasks should be done judiciously, as close to  97% of performance achievable by using the entire FLAN collection can be attained with just a subset of 16 representative tasks alone with $N^{\prime}=200000$.
$N^{\prime}$ FLANv2 Mix SMART
$25000$ 40.05 39.8
$50000$ 41.49 43.03
$100000$ 43.18 44.96
$200000$ 44.73 46.7
$400000$ 46.26 47.65
Table 3: Comparison of SMART with mixture weights of Longpre et al. (2023) for Llama-2-7b. The scores are weighted average of exact match on MMLU and BBH.
4.7 Comparison with FLANv2 Mix
We now compare SMART with FLANv2-mix i.e., by using the mixture weights suggested by Longpre et al. (2023) where the weights 40%, 32%, 20%, 5% and 3% are assigned to the 5 sub mixtures of FLAN 2022 — FLANv1, T0, NIV2, CoT and Dialog respectively and instances are randomly sampled from each sub mixture. Since this method only prescribes weights on sub mixtures and not on individual tasks, we only consider the case where $M^{\prime}=M$. Table 3 contains comparison of SMART vs FLANv2-mix for Llama-2-7b. FLANv2-mix seems to perform better than EPM (Baseline-1) and EM (Baseline-2) in some cases although SMART almost always outperforms FLANv2-mix.
4.8 Ablation Study: Submodular Function
@@
Table 4: Grid Search on submodular functions $f_{1}$ and $f_{2}$ for Llama-2-7b where weighted average of exact match on MMLU and BBH are compared for different choices of $f_{1}$ and $f_{2}$. FL denotes Facility Location, GC denotes Graph Cut and LOGDET denotes Log-Determinant.
@@
Both Stage-1 and Stage-2 of SMART require us to choose submodular functions $f_{1}$ and $f_{2}$ which are best treated as hyperparameters because of uncertainty with respect to which functions are best suited for instruction tuning (Section 3.3). We conduct a grid search on $f_{1}$ and $f_{2}$ using three functions from Table 1, by setting $M^{\prime}=M$ and varying $N^{\prime}$ in $\{25000,50000,100000,200000\}$. In this work, we use $\lambda=0.4$ for Graph-Cut. The grid search (summarized in Table 4) suggests that Graph Cut is optimal for task subset selection, while Facility Location is best for instance selection when $M^{\prime}=M$. The task scaling curves for each combination of $f_{1}$ and $f_{2}$ (for different $N^{\prime}$s) are present in Appendix B. However, in this section, we highlight the case where $f_{1}$ is Graph Cut and $f_{2}$ is varied. We find that optimal $f_{2}$ in this case also depends on $M^{\prime}$. For instance, in Figure 2 where $N^{\prime}=200000$, Facility Location performs the best at higher $M^{\prime}$, highlighting the importance of representation, while Graph Cut and Log Determinant show better performance at lower $M^{\prime}$, highlighting the importance of diversity. We hypothesize this is because — for higher $M^{\prime}$s, each task on average gets a relatively low budget and the need for representation dominates the need for diversity; however when there is sufficient enough budget for each task, i.e., for lower $M^{\prime}$s, the need for diversity takes over.
@@ 4.9 Ablation Study: LLM
We also test the effectiveness of SMART strategy on two other LLMs - Falcon-7B and Mistral-7B. Table 5 presents a comparison of SMART mixture with the baseline mixtures for Falcon and Mistral when $M^{\prime}=M$. For Falcon, we only report MMLU-zero-shot since Falcon gets an exact match of 0.0 for BBH-zero-shot even after fine-tuning. The task-scaling curves for these models are present in Appendix C.
$N^{\prime}$ Data Mix. Mistral Falcon
$25000$ EPM(Baseline-1) 52.69 23.09
EM(Baseline-2) 51.65 21.1
FLANv2-mix 53.24 26.16
SMART 53.33 22.73
$50000$ EPM(Baseline-1) 53.75 19.86
EM(Baseline-2) 53.16 22.96
FLANv2-mix 54.15 25.66
SMART 53.8 26.31
$100000$ EPM(Baseline-1) 54.24 20.88
EM(Baseline-2) 54.57 24.85
FLANv2-mix 54.14 27.47
SMART 53.75 27.22
$200000$ EPM(Baseline-1) 55.25 24.7
EM(Baseline-2) 54.71 24.39
FLANv2-mix 55.27 27.63
SMART 55.4 30.96
Table 5: Comparison of SMART with baselines for Mistral and Falcon. The scores correspond to weighted averages of exact match on MMLU and BBH.